OpenStudy (anonymous):

Determine which two values the following expression is between. 4pi over 5 3 and 4 2 and 3 4 and 5 1 and 2

OpenStudy (anonymous):

(4*pi)/ 5 =(4 * 3.1416) / 5 = 2.51328 The answer is between 2 and 3.